the attitudes towards living together have changed quite a bit in the past decades. the same question was asked in 1994, and 587 out of 1422 respondent saying they agreed or strongly agreed that it is alright for a couple to live together without intending to get married. The test statistic to determine if more than 40% of americans felt that way in 1994 was z=0.99.

a. the p-value for this test is:

b. what is the value of z needed to make a 91% confidential interval
Answer #1



n = 1422
P = X / n = 0.4128


Test Statistic :-


Z = 0.99

P value = P ( Z > 0.99 ) = 1 - P ( Z < 0.99 ) = 0.1611

Critical value for 91% confidence interval